A boat takes 3 hours to go 42 miles downstream with the current. It takes 4.2 hours to go 42 miles upstream against the current. What is the speed of the current? What is the speed of the boat?
Answer by lwsshak3(11628)

A boat takes 3 hours to go 42 miles downstream with the current. It takes 4.2 hours to go 42 miles upstream against the current. What is the speed of the current? What is the speed of the boat?
**
let x=speed of the boat
let c=speed of the current
net speed with current=x+c
net speed against current=x-c
speed=distance/travel time
..
(x+c)=42/3=14
(x-c)=42/4.2=10
..
x+c=14
x-c=10
add, eliminating c
2x=24
x=12
c=14-x=14-12=2
ans:
speed of the boat=12 mph
speed of the current=2 mph